RhondaSommer (rhondasommer):

Given 🔺ABC with a = 9 angle A = 80⁰ and angle B = 35⁰, find b. Round the sine value to the nearest thousandth and the answer to the nearest hundredth. 5.24 6.53 8.28 67.00

RhondaSommer (rhondasommer):

@sweetburger

RhondaSommer (rhondasommer):

\[sin(80)/9=sin(35)/b\]

RhondaSommer (rhondasommer):

right?

OpenStudy (sweetburger):

I was writing it out and you beat me to it.

OpenStudy (sweetburger):

yes that is the proper setup

RhondaSommer (rhondasommer):

so 0.574*9=0.985*b 5.166=0.985b b=5.24?

RhondaSommer (rhondasommer):

@sweetburger

OpenStudy (sweetburger):

yep looks good